OpenStudy (anonymous):

Simplify. Square root sign 48 plz help

OpenStudy (anonymous):

√48 = √(16*3) = 4√3

OpenStudy (lgbasallote):

\[\sqrt{48} = \sqrt{16 \times 3}\] 16 is a perfect square

OpenStudy (anonymous):

So the answer will be 4 square root 3?

OpenStudy (anonymous):

Yes.

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

Thank you